# Distribution-free data-driven smooth tests without $\chi^{2}$

### v2

@@ 6 Summary and discussion
This article introduces a novel framework for constructing asymptotically distribution-free smooth tests that do not rely on the usual $\chi^{2}$ approximation. It is shown that, even when the parameters are estimated, the asymptotic null distribution provides a reasonable approximation when the sample size is only moderately large.
The asymptotic distribution-free property is achieved by relying on the K2 transform. The latter consists of a change of variable in functional space, which enables the construction of an empirical process with a standard asymptotic null distribution. In the context of smooth tests, such a transformation is especially valuable in that it yields a new family of orthonormal bases in $L^{2}(G_{\beta})$ such that, even when the parameters are estimated, test statistics defined by functions in these bases are asymptotically distribution-free.
The projected bootstrap is also discussed as a computationally efficient alternative to the classical parametric bootstrap. In particular, the projection structure induced by parameter estimation allows us to simulate the null distribution of the test statistics of interest without re-estimating the model parameters at each bootstrap replicate. Simulation experiments show that the computational gain attained by projected bootstrap can be substantial compared to the parametric bootstrap, especially when the estimation of the parameters is CPU-intensive.
While the present manuscript focuses on the univariate setting, the proposed framework can be easily adapted to test multivariate parametric models. In particular, when $G_{\bm{\beta}}$ is $D$-dimensional distribution, the reference distribution, $F_{\bm{\gamma}}$ could be chosen to be a product of $D$ univariate distributions $F_{\bm{\gamma},d}$, $d=1,\dots,D$. The $K2$ orthonormal basis in $L^{2}(G_{\bm{\beta}})$ can then be constructed by applying the K2 transform to a tensor product of bases in $L^{2}(F_{\gamma,d})$ (e.g., Algeri, (2021)).
Extensions to smooth tests for regression (Rayner et al.,, 2022) are also possible. In this case, instead of relying on the classical empirical process for i.i.d. data to express the statistics of interest as functionals from it (see Section 2), the random measure to be used is the weighted empirical process proposed in Khmaladze, (2017).
@@ 7 Acknowledgments
The authors thank Charlie Geyer for the valuable feedback.
@@ B Required Orthogonality Conditions of $\widetilde{c}_{\lambda_{k}}$
$\left\langle\widetilde{c}_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=0,$
$\left\langle\widetilde{c}_{\lambda_{3}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},c_{\lambda_{1}}\right\rangle_{G_{\bm{\beta}}}=0,$
$\left\langle\widetilde{c}_{\lambda_{3}},b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U^{2}_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{1}}\right\rangle_{G_{\bm{\beta}}}=0;$
and this can proceed up to $\widetilde{c}_{\lambda_{p}}$ by induction.

### v3

@@ 6 Summary and discussion
This article introduces a new class of smooth test statistics whose asymptotic distribution-freeness is achieved by relying on the K2 transform. The latter consists of a change of variable in functional space, which enables the construction of an empirical process with a standard asymptotic null distribution. In the context of smooth tests, such a transformation is especially valuable in that it yields a new family of orthonormal bases in $L^{2}(G_{\beta})$. Test statistics based on elements of such bases are shown to be asymptotically distribution-free even when the parameters are estimated, and the sample size is only moderately large.
The projected bootstrap is also discussed as a computationally efficient alternative to the classical parametric bootstrap. In particular, the projection structure induced by parameter estimation allows us to simulate the null distribution of the test statistics of interest without re-estimating the model parameters at each bootstrap replicate. Simulation experiments show that the computational gain attained by the projected bootstrap can be substantial compared to the parametric bootstrap, especially when the estimation of the parameters is CPU-intensive.
While the present manuscript focuses on the univariate setting, the proposed framework can be easily adapted to test multivariate parametric models. In particular, when $G_{\bm{\beta}}$ is $D$-dimensional distribution, the reference distribution, $F_{\bm{\gamma}}$ could be chosen to be a product of $D$ univariate distributions $F_{\bm{\gamma},d}$, $d=1,\dots,D$. The K2 orthonormal basis in $L^{2}(G_{\bm{\beta}})$ can then be constructed by applying the K2 transform to a tensor product of bases in $L^{2}(F_{\gamma,d})$ (e.g., Algeri (2021)).
Extensions to smooth tests for regression (Inglot and Ledwina, 2006a; Rayner et al., 2022) are also possible. In this case, instead of relying on the classical empirical process for i.i.d. data to express the statistics of interest as functionals from it (see Section 2), the random measure to be used is the weighted empirical process proposed in Khmaladze (2017).
@@ 7 Acknowledgments
The authors thank Charlie Geyer, Teresa Ledwina, and two anonymous referees for the valuable feedback.
@@ B Required Orthogonality Conditions of $\widetilde{c}_{\lambda_{k}}$
$\begin{split}\left\langle\widetilde{c}_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}&=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}\circ U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}\circ U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=0,\\ \left\langle\widetilde{c}_{\lambda_{3}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}&=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}\circ U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}\circ U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},c_{\lambda_{1}}\right\rangle_{G_{\bm{\beta}}}=0,\\ \left\langle\widetilde{c}_{\lambda_{3}},b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}&=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}\circ U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}\circ U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}\\ &=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}},c_{\lambda_{1}}}\circ U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{1}}\right\rangle_{G_{\bm{\beta}}}=0.\end{split}$
By induction, the same can be done for $\widetilde{c}_{\lambda_{4}},\dots,\widetilde{c}_{\lambda_{p}}$.